OpenStudy (anonymous):

What is the solution of the system of equation? 3x+2y+z=7 5x+5y+4z=3 3x+2y+3z=1

OpenStudy (anonymous):

using linear algebra, put it in rref to get your answer the other method would be to try to cancel one of the variables out so if you multiply the first equation be -1, and then add all 3 of them you will get an equation of one varable, then substitute it in to the other equations and try to cancel out another variable

OpenStudy (anonymous):

Can you help me step-by-step?

OpenStudy (anonymous):

yea sure first multiply the first equation by -1 and then add all 3 equations together

OpenStudy (anonymous):

can you do it for me? need to make sure you understand what im trying to say

OpenStudy (anonymous):

So 3(-1)+2(-1)+(-1)=7

OpenStudy (anonymous):

when you multiply the equation by -1, you need to multiply both sides of the equation \[-1(3x+2y+z) = 7(-1)\] simplifies to \[-3x - 2y\]

OpenStudy (anonymous):

sorry this is the simiplied form the rest got cutoff for some reason \[-3x - 2y -z = -7 \]

OpenStudy (anonymous):

oh, okay, that makes sense

OpenStudy (anonymous):

now that you multiply both sides of the first equation by -1 you want to add all 3 equations together 

OpenStudy (anonymous):

lol this might be a bit of work to do unless i show you how to do it through linear algebra just solve for y in this new equation \[y = \frac{ -3 }{ 5 } - \frac{ 5 }{ 3 }x \] plug it in the first equation

OpenStudy (anonymous):

I'm not sure how to solve that...

OpenStudy (anonymous):

yea.. sorry i dont think i can do it the usual way without linear algebra.. i can show you using matrices if you like?